Grade 9Algebra

If a+b is equals -1,find the value of a cube +b cube+3ab

Navedita sinha
9 Years agoGrade 9

1 Answer

Vikas TU
9 Years ago
Given that a+b = -1    [Cubing on both sides]
(a+b) 3 = (-1)3
a3 + b3 +3ab(a+b) = -1
a3 + b3 +3ab(-1) = -1
a3 + b3 -3ab= -1                    [Adding 6ab on both sides]
a3 + b3 +3ab= 6ab – 1